# Grounding Foundation Models through Federated Transfer Learning: A General Framework

### v10

@@ 1. Introduction
While FMs have shown great success in various general natural language processing (OpenAI, 2023; Liu et al., 2023c; Kamalloo et al., 2023) and computer vision tasks (Kirillov et al., 2023; Rombach et al., 2022), FMs have limitations that prevent them from being adopted to domain-specific applications: (1) FMs typically are pre-trained on publicly available datasets and lack industry-level and domain-specific knowledge, implying the deficiency of applying FMs to applications of various industrialized domains. (2) FMs are trained only up to a certain point in time, and thus, their knowledge is often outdated, thereby making it challenging to apply FMs to scenarios that require real-time updating of knowledge, e.g., in Newspaper publications. To make FMs really excel in addressing challenging domain-specific problems, we need to ground FMs by integrating them with industry-level and domain-specific knowledge. Nevertheless, some powerful FMs are closed-source (e.g., GPT-4 and PaLM). Valuable domain knowledge is often distributed across different enterprises, organizations, or edge devices, stored in their local data warehouses, or encoded in private machine learning models. Consequently, grounding FMs requires transferring knowledge between FM providers and domain knowledge holders. We refer to the former as the ”server” and the latter as the ”client,” as depicted in Figure 1. In addition, domain knowledge often encompasses sensitive and confidential information, and companies or edge devices may have limited computational resources. These circumstances present challenges for grounding FMs:
@@
Federated Transfer Learning (FTL) provides promising solutions to address these obstacles for grounding FMs. We refer to the concept of grounding FMs through FTL techniques as FTL-FM. Specifically, FTL-FM provides servers of FMs and clients of DMs with privacy and security defense tools (Li et al., 2023c), federated learning algorithms (Yang et al., 2019; McMahan et al., 2017), and transfer learning approaches (Pan and Yang, 2010) to collaboratively adapt FMs to domain-specific models or augment FMs with domain-specific knowledge. In recent years, a considerable amount of methods across the spectrum of FTL-FM have been proposed to ground FMs. Subsequently, several survey and vision papers were proposed to overview these works, summarize application scenarios, and discuss potential challenges and solutions (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a; Zhuang et al., 2023; Chen et al., 2023). However, these works either focus on specific techniques of customizing FMs without considering distributed scenarios of knowledge transfer (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a) or provide high-level overview of federated LLM scenarios, challenges, and potential solutions (Zhuang et al., 2023; Chen et al., 2023) without formulating FTL-FM problems and systematically reviewing relevant FTL-FM works.
@@
  • A detailed taxonomy to categorize state-of-the-art FTL-FM works. According to the FTL-FM framework, we raise five specific research issues, including what knowledge to transfer, how to transfer the knowledge, what the threat model is, what information to protect, and how to protect the information (see Section 4). Based on these research issues, we construct a taxonomy to categorize state-of-the-art FTL-FM works (see Table 2). The FTL-FM framework, together with this taxonomy, can serve as guidance for future FTL-FM works.
@@
The rest of the paper is organized as follows. Section 2 overviews related works. Section 3 defines a general framework of FTL-FM and Section 4 presents a taxonomy to categorize state-of-the-art FTL-FM works. Moving forward, Section 5, Section 6, and Section 7 comprehensively review FTL-FM works of three settings, respectively. Subsequently, Section 9 and Section 10 examine efficiency-improving and privacy-preserving methods, respectively. Section 11 discusses opportunities and future directions. Finally, Section 12 presents the concluding remarks for this work. The outline of the core sections of this work is pictorially illustrated in Figure 2.
@@ 3. FTL-FM Framework
Adapting an FM typically comprises three phases: pre-training, fine-tuning, and prompting. Pre-training aims to train an FM from scratch based on broad ranges of large-scale datasets so that the FM can be adaptable to a wide range of downstream tasks. Fine-tuning aims to train a pre-trained FM with specific goals. It can be instruction tuning that enables an FM to follow human’s instructions for solving specific tasks (Ouyang et al., 2022), alignment tuning that aligns an FM’s behaviors with human intentions and values (Ji et al., 2023), or domain-specific tuning that augments the FM with domain-specific or industry-level knowledge. Prompting is to query an FM with prompts for solving various tasks.
@@ 3.1. The Definition of FTL-FM Framework
Figure 5. Illustration of the three settings formulated in the FTL-FM framework (see Definition 3.1). The objective of setting 1⃝ is transferring and adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs), the objective of setting 2⃝ is leveraging clients’ domain-specific knowledge to enhance the server’s FM, and the objective of setting 3⃝ is co-optimizing both the server’s FM and clients’ DMs.
@@
Remark 3 ().
[…]
Remark 4 ().
[…]
Remark 5 ().
@@ 3.2. Formulations of Machine Learning Tasks
Various knowledge transfer approaches have been proposed to achieve each of the three objectives defined in the FTL-FM framework (see Definition 3.1). We overview and formulate representative approaches. These formulations are summarized in Table 1. To better illustrate these formulations, we denote the FM hosted by the server as $f_{\mathcal{W}^{S}_{\text{FM}}}(\cdot)$ parameterized by $\mathcal{W}^{S}_{\text{FM}}$, the DM owned by client $k$ as $g_{\mathcal{W}^{C_{k}}}(\cdot)$ parameterized by $\mathcal{W}^{C_{k}}$, the data owned by the server as $\mathcal{D}^{S}=\{x^{S},y^{S}\}$, and the data owned by client $k$ as $\mathcal{D}^{k}=\{x^{C_{k}},y^{C_{k}}\}$.
@@
Objective Formulation of Specific Federated Transfer Learning Loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};\{x^{C_{k}},y^{C_{k}}\})$ Knowledge Transfer Approach Representative Work
@@
To achieve the objective of setting 1⃝, a straightforward way is through domain adaptation: the server compresses the FM into a relatively smaller model that is adapted to the client’s task and sends the distilled model to the downstream client for further fine-tuning  (Mireshghallah et al., 2022; Yu et al., 2022; Wang et al., 2023c). We formulate this approach as follows.
@@
where $\ell^{S}_{\text{KD}}$ is the knowledge distillation loss for distilling knowledge from the teacher model $f_{\mathcal{W}^{S}_{\text{FM}}}$ to the student model $f^{\prime}_{\mathcal{W}_{\text{SM}}^{S}}$, which is used to initialize downstream client $k$’s domain model $g_{\mathcal{W}^{C_{k}}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training $g_{\mathcal{W}^{C_{k}}}$.
@@
To achieve the objective of setting 2⃝, federated offsite tuning (Xiao et al., 2023; Fan et al., 2023; WeBank, [n. d.]; Chua et al., 2023; Khalid et al., 2023) and federated knowledge distillation (Yu et al., 2023b) are two representative approaches. We formulate their objectives as follows:
@@
In federated offsite tuning (Fan et al., 2023; WeBank, [n. d.]; Chua et al., 2023), the server first sends $\phi_{\mathcal{W}^{S}_{a}}$, $\psi_{\mathcal{W}^{S}_{b}}$, and $\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ to all clients before training. Then, each client trains the adaptors $\phi_{\mathcal{W}^{S}_{a}}$ and $\psi_{\mathcal{W}^{S}_{b}}$ with the help of the emulator $\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ based on its local data following Eq.(10). Next, all clients send their adaptors back to the server, which in turn aggregates clients’ adaptors and plugs the aggregated adaptors back to the FM. In federated knowledge distillation (Yu et al., 2023b), the $f_{\mathcal{W}^{S}_{\text{FM}}}$ at the server is fine-tuned by the representations generated by each client’s $g_{\mathcal{W}^{C_{k}}}$ following Eq.(12). In both approaches, $f_{\mathcal{W}^{S}_{\text{FM}}}$ can be optionally fine-tuned by the server’s local data if available.
To achieve the objective of setting 3⃝, the server’s FM and clients’ DMs are co-optimized through sharing knowledge of each other. This co-optimization objective can be formulated as follows.
@@
$+\ell_{\text{FT}}(f_{\mathcal{W}_{\text{FM}}^{S}}(x^{a}),y^{a})+\ell_{\text{FT}}(g_{\mathcal{W}^{\diamond}}(x^{b}),y^{b}).$
@@
where $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training client $k$’s DM $g_{\mathcal{W}^{\diamond}},\diamond\in\{G,C_{k}\}$ based on local data $\{x^{C_{k}},y^{C_{k}}\}$; $g_{\mathcal{W}^{G}}$ is the global model shared by all clients when $\diamond$ is $G$, while $g_{\mathcal{W}^{C_{k}}}$ is the personalized model of client $k$ when $\diamond$ is $C_{k}$; $\ell_{\text{FT}}$ is the loss for fine-tuning $f_{\mathcal{W}^{S}_{\text{FM}}}$ and $g_{\mathcal{W}^{\diamond}}$ using knowledge in the form of $\{x^{a},y^{a}\}$ and $\{x^{b},y^{b}\}$, respectively. The specific forms of $\{x^{a},y^{a}\}$ and $\{x^{b},y^{b}\}$ are algorithm-dependent.
@@
The federated co-optimizing can be implemented in various forms, Fan et al. (2023) proposed FedCoLLM (Fan et al., 2023), a federated mutual knowledge distillation approach that enables $f_{\mathcal{W}^{S}_{\text{FM}}}$ and $g_{\mathcal{W}^{G}}$ to learn from each other through knowledge distillation. In FedCoLLM, $x^{a}=x^{b}=x^{S}$, $y^{a}=g_{W^{G}}(x^{S})$, and $y^{b}=f_{\mathcal{W}^{S}_{\text{FM}}}(x^{S})$ where $x^{S}$ is public data. Deng et al. (2023) proposed CrossLM, which accomplishes the co-optimization of $f_{\mathcal{W}^{S}_{\text{FM}}}$ and $g_{\mathcal{W}^{C_{k}}},k\in\{1,\dots,K\}$ based on a synthetic dataset generated by the server’s FM $f_{\mathcal{W}^{S}_{\text{FM}}}$. In CrossLM, $y^{a}=y^{b}=y^{S}$ and $x^{a}=x^{b}=f_{\mathcal{W}^{S}_{\text{FM}}}(\mathcal{F}(y^{S}))$ where $y^{S}$ is the given labels and $\mathcal{F}(y^{S})$ is the corresponding label-descriptive prompts.
@@
A special case of objective 3⃝, as formulated in Eq.(5), is that clients’ DMs are initialized with FMs and they are trained with Parameter-Efficient Fine-Tuning (PEFT) techniques. We formulate the objective of each client $k$ as follows.
@@
(13) federated PEFT: $\min_{\mathcal{W}^{G}_{a}}\ell_{\text{TA}}^{C_{k}}(\phi_{\mathcal{W}^{G}_{a}}\oplus g_{\mathcal{W}^{C_{k}}_{\text{FM}}}(x^{C_{k}}),y^{C_{k}}).$
@@
where $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is the local FM of client $k$ while $\phi_{\mathcal{W}^{G}_{a}}$ is the global PEFT module parameterized by $\mathcal{W}^{G}_{a}$ (e.g., prompt model, adaptor, and LoRA) and shared by all clients; $\oplus$ denotes an operator that composes $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ with PEFT module $\phi_{\mathcal{W}^{G}_{a}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for optimizing the global PEFT module $\phi_{\mathcal{W}^{G}_{a}}$ based on client $k$’s local data $x^{C_{k}}$ and $y^{C_{k}}$ while $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is keeping frozen.
The research works, FedPETuning (Zhang et al., 2023d), FedPrompt (Zhao et al., 2023b), FedAdapter (Cai et al., 2023b), and FedIT (Zhang et al., 2023c) adhere to the formulation of Eq.(13). FedPrompt, FedAdapter, and FedIT specifically investigate the impact of prompt tuning, adapter, and LoRA under federated learning, respectively. On the other hand, FedPETuning delves into exploring four PEFT techniques, namely adapter tuning, prefix tuning, LoRA, and BitFit, while operating under the FedAvg protocol.

### v11

@@ 1. Introduction
While FMs have shown great success in various natural language processing (OpenAI, 2023; Liu et al., 2023c; Kamalloo et al., 2023) and computer vision tasks (Kirillov et al., 2023; Rombach et al., 2022), FMs have limitations that prevent them from being adopted to domain-specific applications: (1) FMs typically are pre-trained on publicly available datasets and lack industry-level and domain-specific knowledge, implying the deficiency of applying FMs to applications of various industrialized domains. (2) FMs are trained only up to a certain point in time, and thus, their knowledge is often outdated, thereby making it challenging to apply FMs to scenarios that require real-time updating of knowledge, e.g., in Newspaper publications. To address these limitations and unlock the full potential of FMs in addressing complex domain-specific tasks, it is imperative to integrate FMs with domain-specific knowledge in a timely manner.
Nevertheless, some powerful FMs are closed-source (e.g., GPT-4 and PaLM), and valuable domain knowledge is often distributed across different enterprises, organizations, or edge devices, stored in their local data warehouses, or encoded in private machine learning models. Consequently, grounding FMs requires transferring knowledge between FM providers and domain knowledge holders. We refer to the former as the ”server” and the latter as the ”client,” as depicted in Figure 1.
@@
The distributed nature of the foundation models and domain knowledge presents several challenges for grounding FMs:
@@
Federated Transfer Learning (FTL) provides promising solutions to address these obstacles for grounding FMs as it provides servers of FMs and clients of DMs with privacy and security defense tools (Li et al., 2023c), federated learning algorithms (Yang et al., 2019; McMahan et al., 2017), and transfer learning approaches (Pan and Yang, 2010) to collaboratively adapt FMs to domain-specific models or augment FMs with domain-specific knowledge. We coin the concept of grounding FMs through FTL techniques as FTL-FM. In recent years, a considerable amount of methods across the spectrum of FTL-FM have been proposed to ground FMs. Subsequently, several surveys and vision papers were proposed to overview these works, summarize application scenarios, and discuss potential challenges and solutions (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a; Zhuang et al., 2023; Chen et al., 2023). However, these works either focus on specific techniques of customizing FMs without covering the technology landscape related to FTL-FM (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a) or they focus on federated learning but lack a comprehensive review of related FTL-FM approaches (Zhuang et al., 2023; Chen et al., 2023). Moreover, these works lack a unified framework that encompasses and formulates existing FTL-FM approaches.
@@
  • A detailed taxonomy to categorize state-of-the-art FTL-FM works. According to the FTL-FM framework, we raise five specific research issues, including what knowledge to transfer, how to transfer the knowledge, what the threat model is, what information to protect, and how to protect the information (see Section 4). Based on these research issues, we construct a taxonomy to categorize state-of-the-art FTL-FM works (see Table 2). The FTL-FM framework and the taxonomy can serve as guidance for future FTL-FM research.
@@
The rest of the paper is organized as follows. Section 2 overviews related works. Section 3 defines a general framework of FTL-FM, including the formulations of three FTL-FM settings, general machine learning tasks, specific knowledge transfer approaches, and privacy leakage. Section 4 presents a taxonomy to categorize state-of-the-art FTL-FM works. Moving forward, Section 5, Section 6, and Section 7 comprehensively review state-of-the-art works of the three FTL-FM settings, respectively. Section 8 discusses privacy threats and protections during foundation model inference. Subsequently, Section 9 and Section 10 examine efficiency-improving and privacy-preserving methods, respectively. Section 11 discusses opportunities and future directions. Finally, Section 12 presents the concluding remarks for this work. The outline of the core sections of this work is pictorially illustrated in Figure 2.
@@ 3. FTL-FM Framework
Adapting an FM typically comprises three phases: pre-training, fine-tuning, and prompting. Pre-training aims to train an FM from scratch based on broad ranges of large-scale datasets so that the FM can be adaptable to a wide range of downstream tasks. Fine-tuning aims to train a pre-trained FM with specific goals. It can be instruction tuning that enables an FM to follow human instructions for solving specific tasks (Ouyang et al., 2022), alignment tuning that aligns an FM’s behaviors with human intentions and values (Ji et al., 2023), or domain-specific tuning that augments the FM with domain-specific knowledge. Prompting is to query an FM with prompts for solving various tasks.
@@ 3.1. The Definition of FTL-FM Framework
Figure 5. Illustration of the three settings formulated in the FTL-FM framework (see Definition 3.1). The objective of setting ① is transferring and adapting knowledge of the server’s foundation model (FM) to clients’ domain models (DMs), the objective of setting ② is leveraging clients’ domain-specific knowledge to enhance the server’s FM, and the objective of setting ③ is co-optimizing both the server’s FM and clients’ DMs.
@@
Remark 3.
Objectives of settings ① and ② aim to optimize clients’ DMs $\mathcal{W}^{C_{k}},k=1,\dots,K$ and the server’s FM $\mathcal{W}^{S}_{\text{FM}}$, respectively. Toward this end, other models may be trained as part of the optimization process. For instance, Objective ① may entail training $\mathcal{W}^{S}_{\text{FM}}$, while objective ② often involve training $\mathcal{W}^{C_{k}},k=1,\dots,K$. For the sake of clarity, these factors are not included in the formulation.
Remark 4.
A special case of the objective formulated in Eq.(4) is that the server provides no model and acts as an aggregation function. The DMs of clients $1,...,K$ are initialized with FMs. Thus, the co-optimizing FMs and DMs is reduced to horizontal federated learning (McMahan et al., 2017; Tan et al., 2022) where clients collaboratively train a global domain-specific foundation model $\mathcal{W}^{G}_{\text{FM}}$.
(5) $\begin{split}&\min_{\mathcal{W}^{G}_{\text{FM}}}\mathcal{L}(\mathcal{W}^{G}_{\text{FM}})=\sum^{K}_{k=1}p^{C_{k}}\ell^{C_{k}}(\mathcal{W}^{G}_{\text{FM}})+\eta\epsilon_{p},\\ &\text{subject to}\quad\epsilon_{p}\leq\xi,\\ &\text{where }\,\,\ell^{C_{k}}(\mathcal{W}^{G}_{\text{FM}})=\mathbb{E}_{d\sim\mathcal{D}^{C_{k}}}[\ell^{C_{k}}(\mathcal{W}^{G}_{\text{FM}};d)],k\in\{1,...,K\}\end{split}$
However, since FMs are expensive to train or clients may initialize their DMs with different FMs, clients typically train a global proxy model or a global Parameter-Efficient Fine-Tuning (PEFT) module $\mathcal{W}^{G}_{\text{a}}$ (e.g., adapter and LoRA) by minimizing their local loss $\ell^{C_{k}}(\mathcal{W}^{G}_{\text{a}},\mathcal{W}^{C_{k}}_{\text{FM}})$ while keeping their local FMs $\mathcal{W}^{C_{k}}_{\text{FM}}$ frozen.
(6) $\begin{split}&\min_{\mathcal{W}^{G}_{\text{a}}}\mathcal{L}(\mathcal{W}^{G}_{\text{a}},\mathcal{W}^{C_{1}}_{\text{FM}},\dots,\mathcal{W}^{C_{K}}_{\text{FM}})=\sum^{K}_{k=1}p^{C_{k}}\ell^{C_{k}}(\mathcal{W}^{G}_{\text{a}},\mathcal{W}^{C_{k}}_{\text{FM}})+\eta\epsilon_{p},\\ &\text{subject to}\quad\epsilon_{p}\leq\xi,\\ &\text{where }\,\,\ell^{C_{k}}(\mathcal{W}^{G}_{\text{a}},\mathcal{W}^{C_{k}}_{\text{FM}})=\mathbb{E}_{d\sim\mathcal{D}^{C_{k}}}[\ell^{C_{k}}(\mathcal{W}^{G}_{\text{a}},\mathcal{W}^{C_{k}}_{\text{FM}};d)],k\in\{1,...,K\}\end{split}$
Remark 5.
@@ 3.2. Formulations of Machine Learning Tasks
Various knowledge transfer approaches have been proposed to achieve each of the three objectives defined in the FTL-FM framework (see Definition 3.1). In this section, we review and formulate representative approaches. These formulations are summarized in Table 1. To better illustrate these formulations, we denote the FM hosted by the server as $f_{\mathcal{W}^{S}_{\text{FM}}}(\cdot)$ parameterized by $\mathcal{W}^{S}_{\text{FM}}$, the DM owned by client $k$ as $g_{\mathcal{W}^{C_{k}}}(\cdot)$ parameterized by $\mathcal{W}^{C_{k}}$, the data owned by the server as $\mathcal{D}^{S}=\{x^{S},y^{S}\}$, and the data owned by client $k$ as $\mathcal{D}^{k}=\{x^{C_{k}},y^{C_{k}}\}$.
@@
Objective Formulation of Specific Federated Transfer Learning Loss $\ell^{C_{k}}$ of each client $k$ Knowledge Transfer Approach Representative Work
@@
$\ell_{\text{TA}}^{C_{k}}(g_{\mathcal{W}^{G}_{\text{FM}}}(x^{C_{k}}),y^{C_{k}})$ where $g_{\mathcal{W}^{G}_{\text{FM}}}$ is the global domain-specific foundation model shared by all clients. federated full-model training or fine-tuning  (Lin et al., 2022; Woisetschläger et al., 2023)
@@
To achieve the objective of setting ①, a straightforward way is through domain adaptation: the server compresses the FM into a relatively smaller model that is adapted to the client’s task and sends the distilled model to the downstream client for further fine-tuning  (Mireshghallah et al., 2022; Yu et al., 2022; Wang et al., 2023c). We formulate this approach as follows.
@@
where $\ell^{S}_{\text{KD}}$ is the knowledge distillation loss for distilling knowledge from the teacher model $f_{\mathcal{W}^{S}_{\text{FM}}}$ to the student model $f^{\prime}_{\mathcal{W}_{\text{SM}}^{S}}$, which is used to initialize downstream client $k$’s domain model $g_{\mathcal{W}^{C_{k}}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss to fine-tune local model $g_{\mathcal{W}^{C_{k}}}$ of client $k$.
@@
To achieve the objective of setting ②, federated offsite tuning (Xiao et al., 2023; Fan et al., 2023; WeBank, [n. d.]; Chua et al., 2023; Khalid et al., 2023) and federated knowledge distillation (Yu et al., 2023b) are two representative approaches. We formulate their objectives as follows:
@@
In federated offsite tuning (Fan et al., 2023; WeBank, [n. d.]; Chua et al., 2023), the server first sends $\phi_{\mathcal{W}^{S}_{a}}$, $\psi_{\mathcal{W}^{S}_{b}}$, and $\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ to all clients before training. Then, each client trains the adaptors $\phi_{\mathcal{W}^{S}_{a}}$ and $\psi_{\mathcal{W}^{S}_{b}}$ with the help of the emulator $\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ based on its local data. Next, all clients send their adaptors back to the server, which in turn aggregates clients’ adaptors and plugs the aggregated adaptors back to the FM. In federated knowledge distillation (Yu et al., 2023b), the $f_{\mathcal{W}^{S}_{\text{FM}}}$ at the server is fine-tuned by the representations generated by each client’s $g_{\mathcal{W}^{C_{k}}}$. In both approaches, $f_{\mathcal{W}^{S}_{\text{FM}}}$ can be optionally fine-tuned by the server’s local data if available.
To achieve the objective of setting ③, the server’s FM and clients’ DMs are co-optimized through sharing knowledge of each other. This specific co-optimization objective of each client $k$ can be formulated as follows.
@@
$+\ell_{\text{FT}}(f_{\mathcal{W}_{\text{FM}}^{S}}(x^{a}),y^{a})+\ell_{\text{FT}}(g_{\mathcal{W}^{\diamond}}(x^{b}),y^{b}),$
@@
where $g_{\mathcal{W}^{G}}$ is the global model shared by all clients when $\diamond$ is $G$, and $g_{\mathcal{W}^{C_{k}}}$ is the personalized model of client $k$ when $\diamond$ is $C_{k}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training client $k$’s DM $g_{\mathcal{W}^{\diamond}},\diamond\in\{G,C_{k}\}$ based on local data $\{x^{C_{k}},y^{C_{k}}\}$; $\ell_{\text{FT}}$ is the loss for fine-tuning $f_{\mathcal{W}^{S}_{\text{FM}}}$ and $g_{\mathcal{W}^{\diamond}}$ using knowledge in the form of $\{x^{a},y^{a}\}$ and $\{x^{b},y^{b}\}$, respectively. The specific forms of $\{x^{a},y^{a}\}$ and $\{x^{b},y^{b}\}$ are algorithm-dependent.
@@
The federated co-optimization can be implemented in various forms. Fan et al. (2023) proposed FedCoLLM (Fan et al., 2023) that accomplishes the co-optimization by distilling knowledge between the server’s FM $f_{\mathcal{W}^{S}_{\text{FM}}}$ and the global DM $g_{\mathcal{W}^{G}}$ aggregated from client’s local DMs. In FedCoLLM, $x^{a}=x^{b}=x^{S}$, $y^{a}=g_{W^{G}}(x^{S})$, and $y^{b}=f_{\mathcal{W}^{S}_{\text{FM}}}(x^{S})$ where $x^{S}$ is public data. Deng et al. (2023) proposed CrossLM that accomplishes the co-optimization of $f_{\mathcal{W}^{S}_{\text{FM}}}$ and $g_{\mathcal{W}^{C_{k}}},k\in\{1,\dots,K\}$ based on a synthetic dataset generated by $f_{\mathcal{W}^{S}_{\text{FM}}}$. In CrossLM, $y^{a}=y^{b}=y^{S}$ and $x^{a}=x^{b}=f_{\mathcal{W}^{S}_{\text{FM}}}(\mathcal{F}(y^{S}))$ where $y^{S}$ is the given labels and $\mathcal{F}(y^{S})$ is the corresponding label-descriptive prompts.
@@
A special case of objective ③ is that the server serves no FM while clients’ DMs are initialized with FMs, and client’s domain-specific FMs are trained with full model fine-tuning (see Eq.(5)) or parameter-efficient fine-tuning techniques (see Eq.(6)). We formulate the specific objectives of each client $k$ as follows.
(14) federated full-model training or fine-tuning: $\min_{\mathcal{W}^{G}_{\text{FM}}}\ell_{\text{TA}}^{C_{k}}(g_{\mathcal{W}^{G}_{\text{FM}}}(x^{C_{k}}),y^{C_{k}}).$
@@
(15) federated parameter-efficient fine-tuning: $\min_{\mathcal{W}^{G}_{a}}\ell_{\text{TA}}^{C_{k}}(\phi_{\mathcal{W}^{G}_{a}}\oplus g_{\mathcal{W}^{C_{k}}_{\text{FM}}}(x^{C_{k}}),y^{C_{k}}).$
@@
where $g_{\mathcal{W}^{G}_{\text{FM}}}$ is the global FM, $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is the local FM of client $k$, and $\phi_{\mathcal{W}^{G}_{a}}$ is the global PEFT module parameterized by $\mathcal{W}^{G}_{a}$ (e.g., adaptor and LoRA) and shared by all clients; $\oplus$ denotes an operator that composes $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ with PEFT module $\phi_{\mathcal{W}^{G}_{a}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for optimizing the global FM or PEFT module $\phi_{\mathcal{W}^{G}_{a}}$ based on client $k$’s local data $x^{C_{k}}$ and $y^{C_{k}}$.
In federated full-model training or fine-tuning (Lin et al., 2022; Woisetschläger et al., 2023), clients initialize their DMs with FMs having the same architecture and conduct knowledge transfer by sharing their full DMs with the server for aggregation. In federated parameter-efficient fine-tuning (Zhang et al., 2023d; Zhao et al., 2023b; Cai et al., 2023b; Zhang et al., 2023c), clients initialize their DMs with FMs that may have different architectures and conduct knowledge transfer by sharing partial or proxies of their DMs.